OpenStudy (nathanjhw):

Solve the differential equation dy/dx = 2x^3/y with initial condition y(1) = 2.

OpenStudy (nathanjhw):

I believe the answer is 1/2 y^2 = 1/2x^4 + 4

OpenStudy (nathanjhw):

@Loser66

OpenStudy (nathanjhw):

Or it might be (1/2) y^2 = (1/2)x^4 + 3

OpenStudy (anonymous):

i think it's the latter

OpenStudy (nathanjhw):

How do you know if it is +3 or +4?

OpenStudy (anonymous):

checked on wolfram alpha

OpenStudy (anonymous):

i used to do calculations like this but haven't done them in months so i'm a little rusty.

OpenStudy (nathanjhw):

Can you send me the link?

OpenStudy (nathanjhw):

Well sqrt(x^4+3) s not the same as x^4+3

OpenStudy (nathanjhw):

Actually wait. I understand you square both sides.

OpenStudy (anonymous):

it is if you take the square root on both sides, you you y^2 on one side. that cancels each other out doesn't it?

OpenStudy (nathanjhw):

I saw that mistake soon after I made it.

OpenStudy (nathanjhw):

Thank you very much!

OpenStudy (anonymous):

it's ok, mistakes are easily made in math. I do it all the time

OpenStudy (irishboy123):

\[\frac{dy}{dx} = \frac{2x^3}{y}\] \[\int\limits y dy = \int\limits2x^3 dx\] \[\frac {y^2}{2} = \frac{x^4}{2} + C\] y(1) = 2 \[\frac{2^2}{2} = \frac{1^4}{2} + C; C = \frac{3}{2}\] \[y^2 = x^2 + 3\]